# Haitao Xu Corresponding author    Jingru Zhang

### v1

@@ 3.1 The feasibility test
Recall that for each edge $e$ and each $P_{i}\in\mathcal{P}$, the ordered set $X_{i}$ obtained in the preprocessing contains $\mathsf{E}\mathrm{d}(P_{i},x)$’s turning points on $e$, and each point is associated with $\mathsf{E}\mathrm{d}(P_{i},x)$’s function over the interval between it and its next point. Consequently, it takes $O(m)$ time to solve $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)\leq\lambda$ for $x\in e$, and given any point $p$, value $w_{i}\mathsf{E}\mathrm{d}(P_{i},p)$ can be known in $O(\log m)$ time by a binary search on $X_{i}$.
@@
Given a set $E^{\prime}$ of $k^{\prime}$ edges $e_{1},\cdots,e_{k^{\prime}}$ and a set $V^{\prime}$ of $k^{\prime\prime}$ distinct vertices so that $k^{\prime}+k^{\prime\prime}=k$, the local feasibility test is to determine whether there are $k^{\prime}$ points $q_{1},\cdots q_{k^{\prime}}$ so that for each $1\leq s\leq k^{\prime}$, $q_{s}$ is in the interior of $e_{s}$ and $V^{\prime}\cup\{q_{1},\cdots,q_{k^{\prime}}\}$ cover $\mathcal{P}$ under $\lambda$. Similar to the algorithm [7] for $m=1$, we solve the local feasibility test for every $k$-size subset of $E\cup V$ that however, allows for multiple instances for an edge for addressing the situation where an edge may contain multiple centers under $\lambda$.
@@
Due to $|V|=O(|E|)$, the algorithm runs in $O(|E|^{k}(kn+\tau))$ where $\tau$ is the running time for solving the above special case of the local feasibility test.
@@
Combining with Lemma 3, we have the following result for the feasibility test.
@@
For $k=2$, the feasibility of $\lambda$ can be determined in $O(|E|^{2}m^{2}n\log mn)$ time; for $k>2$, the feasibility of $\lambda$ can be determined in $O(\min\{|E|^{k}m^{k}n^{k+1}k\log m,|E|^{k}n^{\frac{k}{2}}m^{\frac{k^{2}}{2}}\}\})$ time.
@@ 3.2 Computing $\lambda^{*}$
Lemma 2 implies that $\lambda^{*}$ is of the form either $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)=w_{j}\mathsf{E}\mathrm{d}(P_{j},x)$ or decided by a breakpoint of $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ for some $1\leq i\leq n$. To find $\lambda^{*}$, we implicitly enumerate this candidate set of $\lambda^{*}$ as the set of $y$-coordinates of intersections between lines containing line segments on graphs of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ so that we can employ the line arrangement search technique [13] to find $\lambda^{*}$ with the assistance of our feasibility test.
Consider the $x,y$-coordinate system where every edge $e$ is a line segment $[0,\overline{e}]$ on the $x$-axis. Consider for each $1\leq i\leq n$ function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ w.r.t. each edge. Let $L$ be the set of lines containing line segments on graphs of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ and vertical lines through $e$’s endpoints on the $x$-axis. Clearly, $\lambda^{*}$ is the $y$-coordinate of the lowest intersection between lines in $L$ with a feasible $y$-coordinate. $|L|=O(|E|mn)$ and $L$ can be obtained in $O(|E|mn)$ time (since $\mathsf{E}\mathrm{d}(P_{i},x)$ for each edge has been determined in the preprocess work).
@@ 4 Solving the one-center problem
Recall that $X_{i}=\{x_{1},\cdots,x_{m_{i}}\}$ is the ordered set of all turning points on $e$ of function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$, which has been determined in the preprocessing work. Let $X=\{1\leq s\leq M\}$ be the union of $X_{i}$ for all $1\leq i\leq n$ and $M=O(mn)$. Clearly, $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ is a line segment over each interval $[x_{s},x_{s+1}]$ with $1\leq s<M$.
@@
Consider the problem of computing the lowest points on the upper envelope of functions $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$’s over interval $[x_{s},x_{s+1}]$ for all $1\leq s<M$ from left to right. Because the upper envelope of lines is geometric dual to the convex (lower) hull of points [8]. This problem can be addressed by applying the dynamic convex-hull maintenance data structure [8] which supports logarithmic-time insertion, deletion, query about the lowest point on the dual upper envelope, and query about the point where a vertical line intersects the dual upper envelope. The details are presented as follows.
@@
Merge sets $X_{i}$ for all $1\leq i\leq n$ into a whole ordered list $X$; remove duplicates from $X$ while compute for each point $x_{s}\in X^{\prime}$ a subset $\mathcal{P}(x_{s})$ of uncertain points so that each $P_{i}\in\mathcal{P}(x_{s})$ has its function $\mathsf{E}\mathrm{d}(P_{i},x)$ turning at $x_{s}$ and it is associated with its function $\mathsf{E}\mathrm{d}(P_{i},x)$ over $[x_{s},x_{s+1}]$. These operations can be carried out in $O(mn\log mn)$ time by sorting sets $X_{i}$ and then looping $X$ to remove duplicates while computing that subset for each point. Note that $x_{1}\in X$ is a vertex of $e$ and the other vertex of $e$ is the last point $x_{M}$ of $X^{\prime}$; additionally, subset $\mathcal{P}(x_{1})=\mathcal{P}$ and $\mathcal{P}(x_{M})=\phi$.
@@
Proceed to find the lowest point on the upper envelope $\phi_{e}(\mathcal{P},x)$ over intervals $[x_{s},x_{s+1}]$ in order. For $x_{1}\in X$, we construct in $O(n\log n)$ time the dynamic convex-hull maintenance data structure $Y$ [8] for $n$ lines generated by extending the line segment of each $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in[x_{1},x_{2}]$. Then, we perform in order the following three queries on $Y$ in $O(\log n)$ time: the lowest-point query and the queries about the two intersections where vertical lines $x=x_{1}$ and $x=x_{2}$ intersect the upper envelope. For our purpose, we keep only the lowest one of the obtained points in slab $x_{1}\leq x\leq x_{2}$.
@@
Starting with $s=2$, we compute the lowest point of $\phi_{e}(\mathcal{P},x)$ over intervals $[x_{s},x_{s+1}]$ with $2\leq s<M$ as follows. Suppose we are about to compute the lowest point of $\phi_{e}(\mathcal{P},x)$ for $x\in[x_{s},x_{s+1}]$. At this point, $Y$ maintains the upper envelope of lines containing line segments of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ over the previous interval $[x_{s-1},x_{s}]$. Recall that subset $\mathcal{P}(x_{s})$ contains all uncertain point whose function $\mathsf{E}\mathrm{d}(P_{i},x)$ turns at $x=x_{s}$. For each $P_{i}\in\mathcal{P}(x_{s})$, we first delete from $Y$ its function over $[x_{s-1},x_{s}]$ in $O(\log n)$ time, and then insert its function over $[x_{s},x_{s+1}]$, which is associated with $P_{i}$ in $\mathcal{P}(x_{s})$. After these $2|\mathcal{P}(x_{s})|$ updates on $Y$, we find the lowest point of $\phi_{e}(\mathcal{P},x)$ over $[x_{s},x_{s+1}]$ as for interval $[x_{1},x_{2}]$ by performing the three queries for interval $[x_{s},x_{s+1}]$. Last, find the lowest point of $\phi_{e}(\mathcal{P},x)$ over $[x_{1},x_{s+1}]$ in $O(1)$ time.
@@
It is easy to see that the processing time for interval $[x_{s},x_{s+1}]$ is dominated by the term $mn\log mn+2|\mathcal{P}(x_{t})|\log n$. Thus, with $O(mn\log m)$ preprocessing work for determining functions $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e$, the lowest point of $\phi_{e}(\mathcal{P},x)$, i.e., the local center of $\mathcal{P}$ on $e$ and its objective value, can be found in $O(mn\log mn+\log n\cdot\sum_{s=1}^{M}|\mathcal{P}(x_{s})|)$ time, which is $O(mn\log mn)$.
@@ 5 Conclusion
In this paper, we study the $k$-center problem with $n$ (weighted) uncertain points on an undirected graph. We give the first exact algorithms for solving this problem respectively under $k=1$, $k=2$, and $k\geq 3$. Due to the NP-hardness of this problem, it is quite interesting to explore the approximation algorithms for solving it and study the problem on special graphs including the cactus graph, the partial $k$-tree, etc.

### v2

@@ 3.1 The feasibility test
Recall that for each edge $e$ and each $P_{i}\in\mathcal{P}$, the ordered set $X_{i}$ obtained in the preprocessing contains $\mathsf{E}\mathrm{d}(P_{i},x)$’s turning points on $e$, and each point in $X_{i}$ is associated with $\mathsf{E}\mathrm{d}(P_{i},x)$’s function over the interval between it and its next point. Consequently, it takes $O(m)$ time to solve $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)\leq\lambda$ for $x\in e$, and given any point $p$, value $w_{i}\mathsf{E}\mathrm{d}(P_{i},p)$ can be known in $O(\log m)$ time by a binary search on $X_{i}$.
@@
Given a set $E^{\prime}$ of $k^{\prime}$ edges $e_{1},\cdots,e_{k^{\prime}}$ and a set $V^{\prime}$ of $k^{\prime\prime}$ distinct vertices so that $k^{\prime}+k^{\prime\prime}=k$, the local feasibility test is to determine whether there are $k^{\prime}$ points $q_{1},\cdots,q_{k^{\prime}}$ so that for each $1\leq s\leq k^{\prime}$, $q_{s}$ is in the interior of $e_{s}$ and $V^{\prime}\cup\{q_{1},\cdots,q_{k^{\prime}}\}$ cover $\mathcal{P}$ under $\lambda$. Similar to the algorithm [7] for $m=1$, we solve the local feasibility test for every $k$-size subset of $E\cup V$ that however, allows for multiple instances for an edge for addressing the situation where an edge may contain multiple centers under $\lambda$.
@@
Due to $|V|=O(|E|)$, the algorithm runs in $O(|E|^{k}(kn+\tau))$, where $\tau$ is the running time for solving the above special case of the local feasibility test.
@@
Combining with Lemma 3, we have the following result for the feasibility test. Notice that for $k>3$, the first algorithm is more efficient when $m\geq n$.
@@
For $k=2$, the feasibility of $\lambda$ can be determined in $O(|E|^{2}m^{2}n\log mn)$ time; for $k\geq 3$, the feasibility of $\lambda$ can be determined in $O(\min\{|E|^{k}m^{k}n^{k+1}k\log m,|E|^{k}m^{\frac{k^{2}}{2}}n^{\frac{k}{2}}\}\})$ time.
@@ 3.2 Computing $\lambda^{*}$
Lemma 2 implies that $\lambda^{*}$ is of the form either $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)=w_{j}\mathsf{E}\mathrm{d}(P_{j},x)$ or decided by a breakpoint of $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ for some $1\leq i\leq n$. To find $\lambda^{*}$, we implicitly enumerate this candidate set of $\lambda^{*}$ as the set of $y$-coordinates of intersections between lines containing line segments on graphs of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$, so that we can employ the line arrangement search technique [13] to find $\lambda^{*}$ with the assistance of our feasibility test.
Consider the $x,y$-coordinate system where every edge $e$ is a line segment $0,\overline{e}]$ on the $x$-axis. Consider for each $1\leq i\leq n$ function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ w.r.t. each edge. Let $L$ be the set of lines containing line segments on graphs of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ and vertical lines through $e$’s endpoints on the $x$-axis. Clearly, $\lambda^{*}$ is the $y$-coordinate of the lowest intersection between lines in $L$ with a feasible $y$-coordinate. $|L|=O(|E|mn)$ and $L$ can be obtained in $O(|E|mn)$ time (since $\mathsf{E}\mathrm{d}(P_{i},x)$ for each edge has been determined in the preprocess work).
@@ 4 Solving the one-center problem
Recall that $X_{i}=\{x_{1},\cdots,x_{m_{i}}\}$ is the ordered set of all turning points of function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ on $e$, which has been determined in the preprocessing work. Let $X=\{1\leq s\leq M\}$ be the union of $X_{i}$ for all $1\leq i\leq n$ and $M=O(mn)$. Clearly, $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ is a line segment over each interval $x_{s},x_{s+1}]$ with $1\leq s<M$.
@@
Consider the problem of computing the lowest points on the upper envelope of functions $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$’s over interval $x_{s},x_{s+1}]$ for all $1\leq s<M$ from left to right. Because the upper envelope of lines is geometric dual to the convex (lower) hull of points [8]. This problem can be addressed by applying the dynamic convex-hull maintenance data structure [8] which supports logarithmic-time insertion, deletion, query about the lowest point on the dual upper envelope, and query about the point where a vertical line intersects the dual upper envelope. The details are presented as follows.
@@
Merge sets $X_{i}$ for all $1\leq i\leq n$ into a whole ordered list $X$; remove duplicates from $X$ while compute for each point $x_{s}\in X$ a subset $\mathcal{P}(x_{s})$ of uncertain points so that each $P_{i}\in\mathcal{P}(x_{s})$ has its function $\mathsf{E}\mathrm{d}(P_{i},x)$ turning at $x_{s}$ and it is associated with its function $\mathsf{E}\mathrm{d}(P_{i},x)$ over $x_{s},x_{s+1}]$. These operations can be carried out in $O(mn\log mn)$ time by sorting sets $X_{i}$ and then looping $X$ to remove duplicates while computing that subset for each point. Note that $x_{1}\in X$ is a vertex of $e$ and the other vertex of $e$ is the last point $x_{M}$ of $X$; additionally, subset $\mathcal{P}(x_{1})=\mathcal{P}$ and $\mathcal{P}(x_{M})=\phi$.
@@
Proceed to find the lowest point on the upper envelope $\phi_{e}(\mathcal{P},x)$ over intervals $x_{s},x_{s+1}]$ in order. For $x_{1}\in X$, we construct in $O(n\log n)$ time the dynamic convex-hull maintenance data structure $Y$ [8] for $n$ lines generated by extending the line segment of each $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in[x_{1},x_{2}]$. Then, we perform in order the following three queries on $Y$ in $O(\log n)$ time: the lowest-point query and the queries about the two intersections where vertical lines $x=x_{1}$ and $x=x_{2}$ intersect the upper envelope. For our purpose, we keep only the lowest one of the obtained points in slab $x_{1}\leq x\leq x_{2}$.
@@
Starting with $s=2$, we compute the lowest point of $\phi_{e}(\mathcal{P},x)$ over intervals $x_{s},x_{s+1}]$ with $2\leq s<M$ as follows. Suppose we are about to compute the lowest point of $\phi_{e}(\mathcal{P},x)$ for $x\in[x_{s},x_{s+1}]$. At this moment, $Y$ maintains the upper envelope of lines containing line segments of functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ over the previous interval $x_{s-1},x_{s}]$. Recall that subset $\mathcal{P}(x_{s})$ contains all uncertain points whose functions $\mathsf{E}\mathrm{d}(P_{i},x)$ turn at $x=x_{s}$. For each $P_{i}\in\mathcal{P}(x_{s})$, we first delete from $Y$ its function over $x_{s-1},x_{s}]$ in $O(\log n)$ time, and then insert its function over $x_{s},x_{s+1}]$, which is associated with $P_{i}$ in $\mathcal{P}(x_{s})$. After these $2|\mathcal{P}(x_{s})|$ updates on $Y$, we find the lowest point of $\phi_{e}(\mathcal{P},x)$ over $x_{s},x_{s+1}]$ as for interval $x_{1},x_{2}]$ by performing the three queries for interval $x_{s},x_{s+1}]$. Last, find the lowest point of $\phi_{e}(\mathcal{P},x)$ over $x_{1},x_{s+1}]$ in $O(1)$ time.
@@
It is easy to see that the processing time for interval $x_{s},x_{s+1}]$ is dominated by the term $(mn\log mn+2|\mathcal{P}(x_{s})|\log n)$. Thus, with $O(mn\log m)$ preprocessing work for determining functions $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e$, the lowest point of $\phi_{e}(\mathcal{P},x)$, i.e., the local center of $\mathcal{P}$ on $e$ and its objective value, can be found in $O(mn\log mn+\log n\cdot\sum_{s=1}^{M}|\mathcal{P}(x_{s})|)$ time, which is $O(mn\log mn)$.
@@ 5 Conclusion
In this paper, we study the $k$-center problem with $n$ (weighted) uncertain points on an undirected simple graph. We give the first exact algorithms for solving this problem respectively under $k=1$, $k=2$, and $k\geq 3$. Our problem generalizes the continuous graph $k$-center problem by assuming that the location of each demand customer follows a discrete distribution, which is quite practical especially for locations with measurement errors or dynamic customers. Since the input of our problem involves the probabilistic points on graphs, the developed algorithms may provide insights for addressing related problems in probabilistic combinatorics. Further, due to the NP-hardness of this problem, it is quite interesting to explore the approximation algorithms for solving it and study the problem on special graphs including the cactus graph, the partial $k$-tree, etc.